A tree is losing its leaves. 2/3 fall off in the first week, 2/3 of those remaining fall off in the second week, and 2/3 of those remaining fall off in the third week. Now there are 37 leaves. How many leaves did the tree start with?
1
Expert's answer
2020-03-02T16:49:40-0500

After first week remains "1-\frac{2}{3}=\frac{1}{3}" leaves

After second week remains "\frac{1}{3}-\frac{1}{3}\frac{2}{3}=\frac{1}{9}" leaves

After third week remains "\frac{1}{9}-\frac{1}{9}\frac{2}{3}=\frac{1}{27}" leaves

Let x be the number of the leaves, the tree starts with.

We have: "\frac{x}{27}=37"

So, the tree starts with "x=27*37=999" leaves.